OpenStudy (anonymous):

Find the degree 3 Taylor polynomial T3(x) of the function f(x)=(7x−1)^4/3 at a=4 I am having a really hard time figuring out these Taylor Polynomials. Please help!

OpenStudy (amistre64):

find 3 or 4 derivatives right?

OpenStudy (anonymous):

I have f'(x)= (28/3)(7x-1)^(1/3) f "(x)= (196/9) (7x-1)^(-2/3) f3(x) = (2744/27) (7x-1)^(-5/3)

OpenStudy (amistre64):

f = (7x−1)^4/3 f' = 4/3 7 (7x−1)^1/3 f'' = 4.1/3.3 7.7 (7x−1)^-2/3 f''' = 4.1.-2/3.3.3 7.7.7 (7x−1)^-5/3

OpenStudy (amistre64):

evaluate those at x=4

OpenStudy (amistre64):

\[T(x)=f(a)+f'(a)(x-a)+\frac1{2!}f''(a)(x-a)^2+\frac1{3!}f''(a)(x-a)^3+...\]

OpenStudy (amistre64):

.... dropped a ' on that last notation

OpenStudy (anonymous):

thank you!! I had just missed my negative on the last derivative

OpenStudy (amistre64):

good job then ;) that hardest thing about calculus is the algebra lol